Senior management is scheduled to review the labor estimate for developing software for X i The project manager requested an analysis of past performance. The assigned analyst compared the actual labor with the initial estimate for the 30 most recent projects. The results, X-actua-initial estimate (in thousands of hours), are as follows 7.2 4.8 9.1 13.2 3.9 6.7 0.1 2.8 10.2 3.8 4.2 16.5 10.9 7.4 5.1 18.1 3.1 9.5 2.5 12.5 -4.2 5.5 0.5 1.9 15.2 14.1 8.3 11.5 11.9 3.lf X has a normal probability distribution, b)Estimate the probability that the current labor estimate will be exceeded. i.e., the project will experience an overrun
Answer #1

here if we assume that it has a normal probability distribution

Here mean of X = 7.19

Standard deviation of X = 5.5933

Now, we have to find that

Pr(X >0 ; 7.19 ; 5.5933) = 1 - Pr(X < 0 ; 7.19 ; 5.5933)

Z = -7.19/5.5933 = -1.2855

Pr(Overrun) = 1 - Pr(Z < -1.2855) = 1 - 0.01 = 0.90

so it is 90% chance that projecct will experience an overrun.